Two people at different locations at a rock concert listen to the same section of a song and for one the sound level is 80 dB, and for the other it is 95 dB. What is the ratio of the intensities of the sound?
Answer #1

the ratio is = 10*log I1/I2

                   = 10* log(80/95)

                   = 10*8.4= 84

The rock concert is 84 times more intense